OpenStudy (anonymous):

Solve by using elimination y=2x-1 2x-y-1=0

OpenStudy (anonymous):

They are the same equations!!!!

OpenStudy (anonymous):

look if you move the x and y to the lefthand side and the constant to the right hand side then you will see that they are the same equations

OpenStudy (anonymous):

y = 2x - 1 2x - y - 1 = 0 y = 2x - 1 -y = -2x + 1 y = 2x - 1 y = 2x - 1 y will vary according to the value of x ~ without further information they could be anything. ANSWER this system has infinitely many solutions

OpenStudy (anonymous):

Solve y=2x-1 2x-y-1=0